I need help!!! I cannot solve the following questions and my summer school teacher is not helpful (I cant understand a word he says).
1. Simplify 2xsquared+3x-(4xsquared-6x)+(9-2x)
2. solve 5y+2=2(y-6)

Answer by checkley71(8403)

2x~2+3x-4x~2+6x+9=0 or -2x~2+7x+9=0 or 2x~2-7x-9=0 or (2x-9)(x+1)=0 or
2x-9=0 or 2x=9 or x=-9/2 or x=-4.5 & x+1=0 or x=-1